OpenStudy (anonymous):

Am I right on this?... To find the distance from two points in a graph I use the x1,y1 and x2,y2 formula?

jimthompson5910 (jim_thompson5910):

I think you are referring to this formula? \[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\]

OpenStudy (anonymous):

but it looks like this...

jimthompson5910 (jim_thompson5910):

ok you can use that formula above, or you can simply count out the spaces between the two points you can only count out the spaces if a) the points have the same x coordinate (one point is directly above the other) OR b)the points have the same y coordinate (the two points are on the same horizontal level)

OpenStudy (anonymous):

(-7,5) (-7,-3) the points do have the same x.

jimthompson5910 (jim_thompson5910):

so you can count out the spaces between the points if you don't want to count out the spaces, you can subtract the coodinates that aren't the same (in this case, the y coordinates) take the absolute value to ensure the result is positive |5 - (-3)| = ??

OpenStudy (anonymous):

(0,2) ??

OpenStudy (anonymous):

Am i right?

jimthompson5910 (jim_thompson5910):

no, the distance is going to be a single number

OpenStudy (anonymous):

so 2

jimthompson5910 (jim_thompson5910):

5 - (-3) is not equal to 2

OpenStudy (anonymous):

i mean 7

jimthompson5910 (jim_thompson5910):

the two negatives make a positive 5 - (-3) = 5+3 = 8

jimthompson5910 (jim_thompson5910):

count it out: there are 8 spaces between the two points

OpenStudy (anonymous):

i am horrible at math today.

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

you're welcome